A 76.00

pound flask of mercury costs $158.50.

The density of mercury is 13.534 g/cm

3

.

Find the price of one cubic inch of mercury by calculating intermediate values.

What is the price of one pound of mercury?

$



What is the price of 1 g of mercury?

$



What is the price of 1 cm3 of mercury?

$



What is the price of 1 in3 of mercury?


1
Expert's answer
2021-02-04T06:53:42-0500

Solution:1)What is the price of one pound of mercury?

"\dfrac{158.50}{76\;lbs}=2.085" $


2)What is the price of 1 g of mercury?

"76\; lbs\;*\;\dfrac{1000g}{2.2046\;lbs}=34473g"


"\dfrac{129}{34473g}=0.0037" $


3)What is the price of 1 cm3 of mercury?

1 cm3 is equivalent to 13.534 g, so that:


"13.534\;g\;*\;0.0037"$ = "0.050"$


4)What is the price of 1 in3 of mercury?

1 cubic inch = 2.543 = 16.387 cm3


"0.050"$ "*\;16.387\;cm^3=0.82"$